OpenStudy (anonymous):

Suppose a sequence, a1 , a2 , a3 , a4 , a5 , ... What is the formula to find the nth term of a quartic sequence?

OpenStudy (amistre64):

well, quartic relates to a 5th degree poly .. how do you define a quartic sequence?

OpenStudy (anonymous):

No, quartic is a 4th degree polynomial.

OpenStudy (anonymous):

@amistre64 A quartic polynimal is 4rth degree

OpenStudy (amistre64):

.... i was close :)

OpenStudy (amistre64):

i spose quintic is 5th?

OpenStudy (anonymous):

@amistre64 correct

OpenStudy (anonymous):

A quartic sequence is given in the form: \(\bf a_n=an^4+bn^3+cn^2+dn+e\) such that the 4rth differences are all the same. @TURITW So pick any 4rth degree polynomial in that form and you'll have the sequence you want.

OpenStudy (anonymous):

? I do not quite understand what you meant.

OpenStudy (amistre64):

would that relate to the difference tiers? first tier is constant, second teir is linear .... 5th tier is 4th degree

OpenStudy (anonymous):

I know the formula for linear sequence Tn = a1 + (a2 - a1)(n - 1)

OpenStudy (amistre64):

a1 a2 a3 a4 a5 a2-a1 a3-a2 a4-a3 a5-a4 a1+(a2-a1)n , n starts at 0 would be a first tier constant difference

OpenStudy (amistre64):

a2-a1 a3-a2 a4-a3 a5-a4 (a3-2a2+a1) a4-2a3+a2 a1 + (a2-a1)n + (a3-2a2+a1) n(n-1)/2! would be a third tier constant difference

OpenStudy (amistre64):

the pattern is looking a little pascally to me

OpenStudy (amistre64):

the 4rth tier might be: ...+(a4-3a3+3a2-a1) n(n-1)(n-2)/3!

OpenStudy (amistre64):

(a4-2a3+a2) - (a3-2a2+a1) a4-2a3+a2 -a3+2a2-a1 --------------- lol, it is .....

OpenStudy (anonymous):

So... Tn = (a4 - 3a3 + 3a2 - a1)n(n - 1)(n - 2) / 3! ?

OpenStudy (amistre64):

has to add in the other tiers

OpenStudy (amistre64):

\[Tier~1: a_1~/0!\] \[Tier~2: (a_2-a_1)~n/1!\] \[Tier~3: (a_3-2a_2+a_1)~n(n-1)/2!\] \[Tier~4: (a_4-3a_3+3a_2-a_1)~n(n-1)(n-2)/3!\] \[Tier~5: (a_5-4a_4+6a_3-4a_2+a_1)~n(n-1)(n-2)(n-3)/4!\] add them up and let n start at 0

OpenStudy (amistre64):

if you want n to start at 1, then just subtract 1 from all the ns

OpenStudy (anonymous):

I see, seems reasonable. Thanks.

OpenStudy (amistre64):

youre welcome

OpenStudy (anonymous):

Here I will give you an example...When I say fourth differences, this is what I mean: Let's say my quartic sequence is given by: \(\bf t_n=x^4\). Now let's list the first few terms:\[\bf {t_n}= \left\{ 0,1, 16,81,256,625...,n^4 \right\}\]Now observe as I take the first difference, i.e. the difference between successive terms such as \(\bf t_2-t_1,t_3-t_2...\):\[\bf f_{n}=\left\{ 1,15,65,175,369... \right\} \]Now I will take the second differences, i.e. \(\bf f_2-f_1,f_3-f_2...\):\[\bf s_n=\left\{ 14,50,110,194... \right\}\]Now the third differences:\[\bf h_n=\left\{ 36,60,84... \right\}\]And finally the fourth differences:\[\bf r_n=\left\{ 24,24,24... \right\}\]That's how it works. For a quartic polynomial/sequence, the 4rth differences are the same, for a quadratic the second differences and for the a linear (arithmetic) sequence, the first differences are the same. @TURITW

OpenStudy (anonymous):

Oh, but what I wanted is the formula to determine the nth term of the sequence. Anyways, thanks

OpenStudy (anonymous):

@TURITW I already gave you that...you can choose any quartic polynomial such as \(\bf n^4\) and that will give your nth term..

OpenStudy (amistre64):

the idea is not to determine some specific formula for a specific sequence; but rather to generalize a formula to develop if for any sequence